# Understanding the Local Geometry of Generative Model Manifolds

### v1

@@ 2.2.1 Local complexity, $\delta$
Here, $\mathbf{B}$ is an orthonormal matrix of size $P\times E$ with $P\leq E$, ||.$||.||_{1}$ is the $\ell_{1}$ norm operator and $r$ is a radius parameter denoting the size of the locality to compute $\delta$ for. The sum over regions $\omega\in V_{\bm{z}}$ requires computing $\Omega\cap V_{\bm{z}}$ which can be computationally intractable for high dimensions. A proxy for computing the partition for $V_{\bm{z}}$ with small $r$ is counting the number of non-linearities within $V_{\bm{z}}$, since for small $r$, the one can assume that the non-linearities do not fold inside $V_{\bm{z}}$, therefore providing an upper bound on the number of regions according to Zaslavsky’s Theorem [26]. To compute local complexity, we use the method described in [13] to estimate the number of knots intersecting $V_{\bm{z}}$.
Local complexity of CPWL neural networks have previously been connected to the expressivity of Deep Neural Networks [8, 21]. From a geometric standpoint local complexity tells us the smoothness of the learned function for a given locality. While local complexity is defined in terms of the neighborhood of a latent vector $\bm{z}$, in this paper we use it interchangeably with the local complexity of the data manifold around point $\bm{x}=\mathcal{G}(\bm{z})$.
2.2.2 Local scaling, $\psi$
Definition 2.
For a CPWL manifold produced by generator $\mathcal{G}$, the local scaling $\psi_{\omega}$ is constant for every region $\omega$ denoting logarithm of the scaling induced by the affine slope $\bm{A}_{\omega}$ for all vectors $\bm{z}\in\omega$. Local scaling for $\omega$ can therefore be expressed as
Suppose $\mathcal{G}$ has a uniform latent distribution, meaning every region $\omega$ has a uniform probability density function at the input. Under an injectivity assumption for an input space region $\omega$ and $S=\{\bm{A}_{\omega}\bm{z}+\bm{b}_{\omega}\forall\bm{z}\in\omega\}$ the output density on $S$, $p_{S}(\bm{x})\propto\frac{1}{e^{\psi_{\omega}}}$. Therefore for an injective mapping, local scaling can be considered an un-normalized likelihood measure, where higher local scaling for any given $z\in\omega$ signals lower posterior density. For two regions $\omega\in\Omega$ and $\omega^{\prime}\in\Omega$, the difference in uncertainty can therefore be written as:
$H_{\omega}-H_{\omega^{\prime}}=\psi_{\omega}-\psi_{\omega^{\prime}},$ (5)
2.2.3 Local rank, $\nu$.
Definition 3. For a CPWL manifold produced by generator $\mathcal{G}$, local rank $\nu_{\omega}$ is the smooth rank of the the per-region affine slope $\bm{A}_{\omega}$ and can be expressed as:
$\nu_{\omega}$ $=\exp\left(-\sum_{i}^{k}\alpha_{i}\log(\alpha_{i})\right)$
where $\text{where }\alpha_{i}$ $=\frac{\sigma_{i}}{\sum_{i}^{k}\sigma_{i}}+\epsilon.$
Here, $\{\sigma_{i}\}_{i=0}^{i=k}$ are non-zero singular values of $\bm{A}_{\omega}$ and $\epsilon=10^{-30}$ is a constant. The local rank $\nu_{\omega}$ therefore denotes the dimensionality of the tangent space on the data manifold.
$t=0.88T$
$t=0.66T$
$t=0.44T$
$t=0.22T$
$t=0$
$z^{t}$
$\psi^{t}$
$\delta^{t}$
$\nu^{t}$
Figure 3: Local geometric descriptors $\{\psi^{t},\delta^{t},\nu^{t}\}$ computed over the input domain of a 2D toy diffusion model producing distribution $z^{t}$ via reverse diffusion at noise level $t$, for an initial $z^{T}=\mathcal{N}(\mathbf{0},\mathbf{I})$. For any input vector, descriptors are computed by conditioning the diffusion model on $t$. We see that with decreasing $t$, local scaling $\psi$ and local rank $\nu^{t}$ decreases and local complexity $\delta^{t}$ increases in the proximity of the support of $z^{0}$, i.e., the learned manifold by the diffusion model. Therefore, at any $t$ during the reverse diffusion process, we can use the local descriptors computed for lower noise scales, for example $t=0.22T$, to guide generation towards or away from $z^{0}$.
2.3 Example: Toy generator trained on a $f:\mathbb{R}^{2}\rightarrow\mathbb{R}^{3}$ task
Setup. We train a toy generator $\mathcal{G}:{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}^{3}$ with depth 3 and width 20, to map the 2-dimensional latent domain $[-10,10]^{2}$ onto a toy 2-manifold in a 3-dimensional output space. We train the generator with a regression loss on a ground truth manifold defined as a mixture of five gaussian functions. In Fig. 2-left, we present analytically computed visualization [10] of the piece-wise linear manifold learned by the generator as well as the latent space partition $\Omega$ represented by dark lines. Every black line represents a non-linearity of the function that folds/bends the latent space while going from ${\mathbb{R}}^{2}$ to ${\mathbb{R}}^{3}$. Therefore, the black lines are knots of the continuous piecewise affine spline generator. Each convex region $\omega$ formed by the intersection of the black lines, is mapped to $\text{Im}(\mathcal{G})$ via per region parameters as described in Equation 1. Each region in the input and output partition is colored by $\psi_{\omega}$. We also use a kernel density estimator (KDE) to estimate the density of generated samples (right) on the data manifold for a uniform latent distribution (middle-right), and color samples with the estimated density.
Observations. After training the generator, the pre-activation zero-level sets (Fig. 2) of neurons are positioned where changes to the slope of the manifold is required. The density of linear regions, i.e., local complexity is higher in the center and lower towards the edges. $\bm{A}_{\omega}$ is full-rank $\forall\omega\in\Omega$. Contrasting the KDE estimate of density and local scaling, we can see that for higher $\psi$ (blue hue) we have lower estimated density (blue hue) and vice-versa.
3 Exploring Generative Model Manifolds using Descriptors
In this Section we will be exploring the geometry of the data manifolds learned by various generative models, e.g., denoising diffusion probabilistic models, latent diffusion models like Stable Diffusion [22].
3.1 DDPM trained on toy $f:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ generation task.
Setup. We train a denoising diffusion probabilistic model [9] on a toy dataset [^1] https://jumpingrivers.github.io/datasauRus to visualize how the local geometry of the learned manifold varies with 1) de-noising steps/noise levels 2) gradient descent steps. In Fig. 3 we present the local complexity $\delta^{t}$, local scaling $\psi^{t}$ and local rank $\nu^{t}$ at timesteps $t\in\{6,17,28,39,50\}$. We consider a grid of points in the 2D data space and compute the local descriptors at each point. Note that we drop the $\omega$ notation as we don’t compute $\Omega$ and no longer compute the scaling and rank in a region-wise fashion. To obtain $\bm{A}_{\omega}$ at any vector $z\in\omega$ where $\omega$ is unknown, we can take the input-output jacobian at $z$ (see Eq. 1). We add a superscript $t$ to denote the descriptors being computed for the learned output manifold at denoising step $t$. $\psi^{t}$ and $\nu^{t}$ therefore denotes the local scaling and rank by the mapping from $t$ to $t+1$.
Observations. We see that the local complexity is higher around the support of $z^{t}$ for all noise levels $t$ and both training steps shown in Fig. 3 and Fig. 19. Both local scaling and rank are lower around the data manifold until the last few diffusion timesteps, when the variance of $\psi$ and $\nu$ diminishes. The qualitative results suggest that given two diffusion models (here one is an early checkpoint and the other is a late checkpoint), local geometry around the training data for the better trained model would have higher $\delta^{t}$, and lower $\psi^{t},\mu^{t}$.
3.2 VAE training dynamics for MNIST
$\delta$
$\psi$
optimization steps
[…]
Observations. By increasing the noise we control the puffiness of the target manifold. We observe that as the noise standard deviation is increased there is 1) increase in $\delta$ indicating the manifold becomes less smooth 2) decrease in local scaling indicating that the uncertainty decreases. We can also observe an initial dip in both local complexity and local scaling. This is similar to what was observed for discriminative models in [13] where a double descent behavior was reported in the local complexity training dynamics of classification models. Based on these results, contrary to the observation in [13], generative models do not have a double descent in local complexity however we do observe a double ascent in local scaling. Our observations suggest that the training dynamics need to be taken into account, when comparing the local manifold geometry between two separately trained models.
3.3 Pre-trained Latent Diffusion Foundation Model
In this section we explore the local geometry of a pre-trained latent diffusion model namely Stable Diffusion v1.4 [22]. Stable Diffusion (SD) comprises of an unconditional vector quantized VAE (VQVAE) and a conditional diffusion model in the latent space of the VQVAE. The diffusion model is trained to map a gaussian distribution to the latent domain of the decoder, therefore we focus our experiments on studying the geometry of the unconditional decoder.
@@ 3.3 Pre-trained Latent Diffusion Foundation Model
Refer to captionIncreasing $\psi$
Figure 6: Local Scaling for encoded ImageNet images. ImageNet images ordered along the columns (from left to right), with increasing local scaling $\psi$ of the Stable Diffusion decoder learned manifold. We observe that ImageNet samples with lower values of $\psi$ contain simpler backgrounds with modal representation of the object category. Conversely for higher $\psi$ we have increasing diversity both in background and foreground features.
3.3.1 Visualizing the local geometry
@@
We use three prompts "a cat", "a dog" and "a fox" to generate three latent vectors using the SD diffusion model. We consider a 2D slice in the latent space, going through the three denoised latents, and compute the local descriptors on a $512\times 512$ uniformly spaced grid. We compute $\delta$ for a $4$-dimensional neighborhood of radius $r=0.00001$. To make $A_{\omega}\mathds{1}_{z\in\omega}$ computation tractable, for any latent vector $z$ we take a $120$-dimensional random orthonormal projection of the SD decoder manifold [3] and compute the input-output jacobian with the projection as the output. We use a JAX implementation of Stable Diffusion on TPU resulting in $3.6s$ required for computing jointly $\delta$, $\psi$, and $\nu$ for one latent vector $z$. In Fig. 5 we present the $\delta,\psi,\nu$, on the 2D slice, along with decoded versions of the three latent vectors, used to compute the 2D slice. In Appendix Figs. 2526212223, and  24 we present generated images from the high or low local descriptor regions from the 2D slice.
@@
Similar to what was observed for a DDPM trained on toy data (see Sec. 3.1), we observe that 1) in the convex hull of the three denoised latents, we have higher complexity indicating its proximity to the SD decoder data manifold 2) lower rank in the convex hull affirms that the convex hull has proximity to the data manifold as well. However we see a sparse collection of latents for which the rank drops further 3) as we move inside the convex hull, local scaling increases. This indicates that a convex combination of "a fox", "a cat" and "a dog" is higher uncertainty w.r.t the SD decoder. However if we move away from the convex hull we see that both $\delta$ and $\psi$ decreases while rank increases. This indicates that as we move away from the data manifold, the local scaling descriptor collapses as such regions are no longer within the support/latent domain of the SD decoder. This is further illustrated through qualitative examples in Appendix Fig. 21 and Fig. 22.
3.3.2 Local geometry of ImageNet
Setup.
Given the observations from the previous sections, we ask the question "What can local geometry of a foundation model say about a dataset?". To find an answer, we consider $20K$ samples from Imagenet with resolution higher or equal to $512\times 512$, encode the samples using the SD encoder, and study the local geometry of the SD decoder manifold for the encoded Imagenet latents. In Fig. 6 we present samples from Imagenet ordered with increasing $\psi$. In Appendix Fig. 14 we present images where the manifold dimensionality is the lowest and highest for the Imagenet samples considered.
Observations.
@@ 3.3.2 Local geometry of ImageNet
In Fig. 6 each column represents a local scaling level set, with increasing values from left to right. Recall that in Eq. 5 we show that increase in local scaling is equivalent to increase in uncertainty. In this figure, we can see that for lower uncertainty images we have more modal features in the images, i.e., the samples have less background elements and are focused on the subject corresponding to the Imagenet class. For higher uncertainty images, we see that images have more outlier characteristics. For images with higher local in Fig. 14, we see that the backgrounds have higher frequency elements compared to lower rank images. For higher rank images, the dimensionality of the manifold is higher locally, therefore allowing more noise dimensions on the manifold. In Appendix. Fig. 15 we present class-wise examples for high and low local scaling images.
Local Scaling, $\psi$
Local Rank, $\nu$
Figure 7: Geometry distinguishes out-of-distribution regions. We encode chest X-ray vs Imagenet images and present here local scaling (left) and local rank (right) distributions computed for the Stable Diffusion decoder. Here chest X-ray images are representative of out-of-distribution samples with respect to the Stable Diffusion learned distribution, and Imagenet images are considered in-distribution. We see that for OOD samples, the expected $\psi$ is higher, showing higher uncertainty under the learned distribution. The local rank is also higher for OOD samples, which correlates with our observations from Fig. 5. These results indicate that the local geometry can be used to characterize samples that are out-of-distribution with respect to a pre-trained generative model.
3.3.3 Detecting Out-of-Distribution Samples
Setup.
To study if the geometric descriptors are discriminative of out-of-distribution samples, we consider evaluating the metrics for $20K$ randomly selected Imagenet samples representing in-domain images for Stable Diffusion and $20K$ X-Ray images from the CheXpert [14] dataset representing out-of-domain images for Stable Diffusion. We present the distributions in Fig. 7.
Observations.
We see that X-ray images have higher uncertainty and higher local rank in expectation compared to Imagenet images. Especially for local rank $\nu$, Chest X-ray images from Chexpert and imagenet image distributions are almost separable. This corroborates with our visualizations from Fig. 5, where away from the convex hull of the denoised latents, we see an increase in $\nu$.
3.3.4 Relationship with Quality, Diversity & Memorization
Quality and diversity of generation for a pre-trained model, is directly connected to the estimation of the target manifold. In this section, we study the relationship between the geometric descriptors and generation quality (good estimation of the manifold), generation diversity (large support of the learned distribution) and memorization (interpolation of the target manifold at a data point).
Setup.
The experiments are organized as such:
$\bullet$ We denoise a fixed set of latent vectors using a) varying guidance scale $\{0,1,3,5,7.5,9\}$ to control quality-diversity trade-off or, b) with memorized/non-memorized prompts by Stable Diffusion. We obtain memorized prompts for SD from [25] and non-memorized prompts from the COCO dataset validation set.
$\bullet$ For a) $50K$ real ImageNet images b) $50K$ synthetic images generated by SD from increasing descriptor level sets we compute the Vendi Score [5] of clip embeddings, the aesthetic score predicted by an aesthetic reward model and a artifact reward model. For the synthetic Imagenet images, we use a guidance scale of $7.5$ and the prompt template "an image of $p$" where $p$ is a randomly class from Imagenet.
Observations.
We see in Fig. 8-bottom that for higher classifier free guidance scales during denoising, uncertainity is lower especially during the final denoising steps. This shows a clear correlation with quality since higher guidance scales result in higher quality images [22]. We also see that for higher guidance scales local rank $\nu$ is $\downarrow$ and local complexity $\delta$ is $\downarrow$ as well. For memorized prompts, in Fig. 8-top, we see that the $\psi$ is significantly lower than non-memorized prompts, indicating very low uncertainity in memorized samples compared to non-memorized samples. We also observe the manifold to be higher dimensional for memorized prompts and smoother.
@@ 3.3.4 Relationship with Quality, Diversity & Memorization
Figure 8: Local geometry of denoising trajectories. Geometric descriptors computed for the VQGAN decoder, during $50$ stable diffusion denoising steps, for (top) 100 COCO and 100 memorized prompts [25] with guidance scale $7.5$ and (bottom) 100 COCO prompts with varying guidance scales. For each prompt or guidance scale, we start from the same seeds. Shaded region represents 95% confidence interval. We see that the local geometry trajectories are discriminative of memorization, as well as increased quality when stronger classifier free guidance is used.
Refer to captionVendi ScoreAesthetic ScoreArtifact ScoreLevel Set Index
Figure 9: Local Geometry level sets for Imagenet images. Vendi scores, Aesthetic scores and Artifact scores computed for 50k generated and real samples of ImageNet from local scaling (top) and local rank (bottom) descriptor level sets of Stable Diffusion. Scores are computed for each level set separately. Synthetic samples are generated using a guidance scale of $7.5$. We see that diversity increases for higher scaling and rank level sets for both real and synthetic data. But the aesthetic/artifact score remains more or less constant for synthetic data, since a high guidance scale is used. This shows that by using local geometry to guide generation, we can increase the diversity of generation while maintaining generation aesthetics.
In Fig.9, we see that with increasing $\psi$ and $\nu$ diversity increases, however for very high $\psi$, i.e., very uncertain images, the diversity collapses. This is due to the higher $\psi$ level sets converging towards the anti-modes of the learned distribution. We see that for real images, aesthetic and artifact scores (higher if without artifacts) get reduced for higher local scaling and rank level sets. However for generated images for higher uncertainty the rewards saturate. This is an indication that even with less perceptual/aesthetic changes, we can generate higher uncertainty images.

### v2

@@ 2.1.3 Local complexity, $\delta$
Here, $\mathbf{B}$ is an orthonormal matrix of size $P\times E$ with $P\leq E$, ||.$||.||_{1}$ is the $\ell_{1}$ norm operator and $r$ is a radius parameter denoting the size of the locality to compute $\delta$ for. Here we consider a $P$ dimensional neighborhood instead of the full dimensionality of the latent space to reduce computational complexity. The sum over regions $\omega\in V_{\bm{z}}$ requires computing $\Omega\cap V_{\bm{z}}$ which can be computationally intractable for high dimensions. A proxy for computing the partition for $V_{\bm{z}}$ with small $r$ is counting the number of non-linearities within $V_{\bm{z}}$, since for small $r$, the one can assume that the non-linearities do not fold inside $V_{\bm{z}}$, therefore providing an upper bound on the number of regions according to Zaslavsky’s Theorem (Zaslavsky, 1975). To compute local complexity, we use the method introduced by Humayun et al. (2024) for general neural networks. We provide in appendix further implementation details and pseudocode.
2.2 Extending beyond Continuous Piecewise-Linear Generators
Computing jacobians for large networks. For any latent vector $\bm{z}\in\omega$, $\bm{A}_{\omega}$ can be obtained by computing the input-output jacobian of the network. Computing the singular values of the full input-output jacobian is significantly expensive for large networks. Therefore, when computing local scaling and rank we obtain singular values via randomized SVD Halko et al. (2011). First we obtain a random projection matrix with orthonormal rows $\mathbf{W}$ with shape $k\times n$ such that $\mathbf{W}\mathbf{W}^{T}=\mathbb{I}_{k}$. Here $n$ is the dimensionality of the outputs generated by the network. We therefore approximate local scaling as:
$\psi_{\omega}^{(trunc)}=\sum_{i=1}^{k}log(\sigma_{i}^{(trunc)})$, where $\sigma_{i}^{(trunc)}$ are the non-zero singular values of $\mathbf{W}\mathbf{A}_{\omega}$.
For any $\omega$, if $\mathbf{W}$ forms a basis for the range of $\mathbf{A_{\omega}}$ then $\sigma_{i}\approx\sigma_{i}^{(trunc)}\forall i={1,2…k}$ [4]. Therefore $\mathbf{WA_{\omega}}$ would provide us a low-rank approximation of $\mathbf{A_{\omega}}$.
In our experiments we have tried two methods to obtain the projection matrix $W$ 1) by obtaining the eigenvectors for the covariance matrix for a set of 50K randomly generated samples. This was suggested in Halko et al. (2011). 2) by performing QR decomposition of a randomly initialized matrix. We see that the performance difference between methods 1) and 2) are negligible therefore consider the cheaper alternative 2) and consider a fixed pre-computed $\mathbf{W}$ with k=120 for all $\mathbf{A_{\omega}}$ in our Stable Diffusion experiments.
Networks with smooth activations. While the descriptors are defined for CPWL mappings, modern generative models employ a mixture of CPWL and non-CPWL operations. For networks with smooth activation functions or non-piecewise-linear non-linearities, our descriptors are first order Taylor approximations. For example, Stable Diffusion employs the GeLU activation function for which we perform the bulk of our experiments in following sections. Smooth activation functions induce a soft VQ partitioning of the latent space compared to the hard VQ partitioning induced by a CPWL map Balestriero & Baraniuk (2018b), retaining much of the local linear structure we expect in CPWL maps. Recent work has also empirically verified the local linearity for a large class of image based diffusion models Chen et al. (2024) suggesting the reliability of first-order approximations.
3 Characterizing the Local Geometry of Pre-Trained Models via Descriptors
In this section, we explore the geometry of pre-trained generative models by characterizing the latent space to output manifold mapping in terms of the local geometric descriptors mentioned in the previous section. We are interested in the following questions: i) How does the on manifold local geometry vary from the off manifold local geometry? ii) How does the local geometry vary across the input domain?
3.1 On and Off Manifold Geometry for Denoising Diffusion Probabilistic Models
$\psi_{t=0.22T}$
$\delta_{t=0.22T}$
$\nu_{t=0.22T}$
$\mathbb{E}_{\mathcal{M}}[\psi_{t}]-\mathbb{E}_{\bar{\mathcal{M}}}[\psi_{t}]$
$\mathbb{E}_{\mathcal{M}}[\delta_{t}]-\mathbb{E}_{\bar{\mathcal{M}}}[\delta_{t}]$
$\mathbb{E}_{\mathcal{M}}[\nu_{t}]-\mathbb{E}_{\bar{\mathcal{M}}}[\nu_{t}]$
Figure 3: Local geometric descriptors computed over the input domain of a DDPM trained on samples from a toy dinosaur manifold $\mathcal{M}\in\mathbb{R}^{2}$, conditioned on $t=0.22T$ (left three columns). Denoising dynamics of local descriptors (right three columns) for different number of training optimization steps. X-axis represents noise level $t=[T,0]$.
Setup. To study the on and off manifold geometry of diffusion models, we train a denoising diffusion probabilistic model (DDPM) (Ho et al., 2020) on a toy dataset [^1] https://jumpingrivers.github.io/datasauRus to visualize how the local geometry varies for 1) different noise levels $t$, and 2) different training iterations. In Fig. 3-heatmaps we present the local complexity $\delta^{t}_{x}$, local scaling $\psi^{t}_{x}$ and local rank $\nu^{t}_{x}$ computed for different input space vectors $x$ using the DDPM conditioned on noise levels $t$. Here $T$ is the highest noise level in the forward diffusion process. We also present the difference between the expected descriptor values on and off the manifold, $\mathbb{E}_{\mathcal{M}}[\Phi]-\mathbb{E}_{\bar{\mathcal{M}}}[\Phi],\forall\Phi\in\{\psi^{t},\delta^{t},\nu^{t}\}$ at different training iterations (right). We consider the set of input vectors within $0.05$ units of the training data as on manifold $\mathcal{M}$ and rest as off the manifold $\bar{\mathcal{M}}$.
Observations. The first observation is that with longer training, the maximum absolute difference between on and off manifold local geometry $\max_{t}\{|\mathbb{E}_{\mathcal{M}}[\Phi]-\mathbb{E}_{\bar{\mathcal{M}}}[\Phi]|\}$ increases. Since with more training we see higher distinction between the on and off manifold geometry, this difference can be an indicator of learning in diffusion models. We see that for well trained models, apart from $t>0.17T$, $\psi^{t}_{x}$ and $\nu^{t}_{x}$ decreases and $\delta^{t}_{x}$ increases with decreasing $t$, $\forall x\in\mathcal{M}$. This means, the likelihood on the manifold increases as noise levels are reduced, the smoothness decreases and the dimensionality of the manifold decreases as well. The quantity $\mathbb{E}_{\mathcal{M}}[\Phi]-\mathbb{E}_{\bar{\mathcal{M}}}[\Phi]$ is also minimized at $t\approx 0.17T$. This indicates that there can exist a noise level $t$ conditioned on which diffusion model local scaling, rank and complexity have the highest distinction geometrically between on and off manifold vectors from the input space. It can allow directly probing which parts of the input space are on the learned manifold to possibly perform one step denoising or propose novel guidance schedules.
@@ 3.1 On and Off Manifold Geometry for Denoising Diffusion Probabilistic Models
3.2 The local geometry of latent diffusion models
In Sec. 3.1, we see that the local geometry in the input domain of a ddpm can be distinctive of its learned manifold. In this section we study the local geometry of the Stable Diffusion (SD) latent space, to explore whether there exists a relationship between the local geometry and the domain of the SD decoder.
@@ 3.2 The local geometry of latent diffusion models
While in Sec. 3.1 we could visualize the whole input domain of the diffusion model, for SD we can only visualize a subspace of the SD latent space. We use three prompts "a cat", "a dog" and "a fox" to generate three latent vectors using the SD diffusion model and consider a 2D slice in the latent space, going through the three denoised latents as our domain to visualize. Note that since this is a 2D subspace of the latent space, we can expect part of it to be in-domain for the SD decoder, whereas part of it would be out-of-domain.
@@
Hum. Pref. Score     Vendi Score
Increasing $\psi$ level sets $\rightarrow$
Figure 5: Local Geometry level sets Imagenet prompts. Vendi diversity scores and RAHF Liang et al. (2024) aesthetic scores computed for images with classifier free guidance (CFG) 7, 5 and 3. Diversity per level set increases and then decrease with increased local scaling. Aesthetic score slightly increases and then decreases as well with increased local scaling.
We observe that 1) In the convex hull of the three denoised latents used as anchors for the 2D subspace being visualized, we have higher complexity, lower rank and higher local scaling. The decoded images from the convex hull may contain artifacts but are legible generations. 2) Local rank does not smoothly vary across the latent space, especially with sharp changes in the local rank within the convex hull of the anchor latent vectors. For the lowest rank regions in the convex hull, decoded images have good fidelity compared to latents with high uncertainty or complexity. 3) If we move away from the convex hull, we see that generated images become more broken and contain heavy artifacts, indicating that such regions are out-of-domain for the SD decoder. However, we see that the local scaling is lower in these regions compared to the convex hull.
4 Secrets that your manifolds hold
Figure 6: Local scaling is sensitive to image corruptions. We use $16$ corruptions from (Hendrycks & Dietterich, 2019) to corrupt $10K$ imagenet images and compute the local geometry of the SD decoder. We see that SD local geometry is sensitive to corruptions, i.e., aesthetic changes to in-domain images (here Imagenet).
Visually Complex images have higher local scaling
@@ 4 Secrets that your manifolds hold
We selected $20K$ samples from Imagenet with resolution higher or equal to $512\times 512$, encode the samples using the SD encoder, and compute the local descriptors for the SD decoder. In Fig. 17 each column represents a local scaling level set, with the $\psi$ for columns increasing from left to right. Recall that in Eq. 3 we show that increase in local scaling is equivalent to increase in uncertainty. In this figure, we can see that for lower uncertainty images we have more modal features in the images, i.e., the samples have less background elements and are focused on the subject corresponding to the Imagenet class.
For higher uncertainty images, we see that images have more outlier characteristics. For images with higher local rank in Fig. 19, we see that the backgrounds have higher frequency elements compared to lower rank images. For higher rank images, the dimensionality of the manifold is higher locally, therefore allowing more noise dimensions on the manifold. These qualitative results provide evidence that the local geometry is indeed sensitive to natural variations of Imagenet images.
Local geometry is sensitive to Image Corruptions
Fig. 6 illustrates the effect of applying 16 different image distortions (originally proposed in (Hendrycks & Dietterich, 2019)) to 10k ImageNet images. We consider Imagenet as in-domain for Stable Diffusion and encode them to the SD latent space to compute the geometric descriptors for the SD decoder. Samples are uniformly distributed over its classes. The plot shows the local scaling distribution at 6 increasing levels of severity $\in\{0,1,2,3,4,5\}$, with zero corresponding to no corruptions applied. We observe that corruptions that are associated with reduction of spectral band, and/or reduction to the color range result in a reduction to the local scaling therefore the negative log-likelihood. We conjecture that this is due to the averaging effect of such distortions which move the corrupted images close to the mean of all images. Conversely, distortions known to be associated with the introduction of high-frequency artifacts are observed to produce an increase in local scaling therefore uncertainty moving the images away from the mean. The results clearly indicate that the local geometry is sensitive to aesthetic changes to images introduced via most of the $16$ corruptions.
@@
Figure 7: Local geometry of denoising trajectories. Geometric descriptors computed for the SD decoder unconditionally, during $50$ stable diffusion denoising steps, for (top) 100 COCO and 100 memorized prompts (Wen et al., 2024) with guidance scale $7.5$ and (bottom) 100 COCO prompts with varying guidance scales. For each prompt or guidance scale, we start from the same seeds. Shaded region represents 95% confidence interval. We see that the local geometry trajectories are discriminative of memorization, as well as increased alignment when stronger classifier free guidance is used.
Memorized and aligned denoising trajectories are locally contracted and relatively smoother.
Classifier-free-guidance is a method for increasing the alignment between the conditioning text-prompt and generated image Rombach et al. (2021). In Fig. 7-bottom, we see that for higher classifier free guidance scales during denoising, the avg. local scaling is lower, avg. local rank is lower and the avg. local complexity is higher. This indicates that for more conditionally aligned images obtained via classifier-free-guidance of the SD diffusion Unet, the decoder uncertainty is also lower especially during the final denoising steps. We also compute the local geometric descriptors for denoising trajectories conditioned on memorized prompts Wen et al. (2024) vs coco captions (Fig.7-top). We see that the mean local geometry is significantly different for denoising trajectories of 100 memorized prompts vs 100 random coco prompts. We also see that for higher guidance scales local rank $\nu$ is $\downarrow$ and local complexity $\delta$ is $\downarrow$ as well.
Connections with Downstream Diversity and Human Preference Scores.
In Fig.8-left we present Vendi score aggregates and in middle and right, we present aesthetic and artifact score (higher is better) aggregates for real and generated Imagenet images with classifier free guidance scales of 7, 5 and 3, sorted in increasing local geometry level sets from left to right. Aesthetic and artifact scores are predicted by a human preference model Liang et al. (2024). We see that for increasing local scaling level sets, we have an increase in the diversity of images per bin. With very high local scaling, we get images from the highest uncertainty modes, i.e., the anti-modes, which result in a drop in the diversity of images. For real images, aesthetic and artifact scores get reduced for higher local scaling non-monotonically. This behavior is expected, as we have discussed before, higher local scaling images have higher uncertainty and higher visual complexity which may result in lower human preference predictions. For images from the highest local scaling bins however, we see an increase in the aesthetic and artifact scores. This is a unique phenomenon showing that for very uncertain images there could be a possible positive human preference. For local rank there is an increase in human preference scores for the high rank bins.